OpenStudy (anonymous):

Find the number of three-digit numbers which when divided by 11 leave a remainder of 9 and when divided by 7 leave a remainder of 2.

OpenStudy (kinggeorge):

Have you heard of the Chinese remainder theorem?

OpenStudy (anonymous):

Do you know that: Dividend = Divisor*Quotient + Remainder Here, we have to find Dividend that is a 3 Digit Number... Getting or not???

OpenStudy (kinggeorge):

I'll assume you haven't. Basically, all you need to know, that there is only one integer out of 77 integers that satisfy the property you want. If you look at \(n=9\), you find that this is the first such integer, since \(9=7*1+2=0*11+9\). Hence, the integers that satisfy the desired property are of the form \(x=a\cdot77+9\). Now we need to find how many are in between 100 and 1000.

OpenStudy (kinggeorge):

How do we do this? Simple. We find the number between 0 and 1000, and subtract off those in between 0 and 100. Can you solve for \(a\) and \(m\) here? \[77\cdot a+m=1000\]

OpenStudy (anonymous):

Don't you think @KingGeorge you are going to very difficult level to solve this question..??

OpenStudy (aravindg):

@PUNYASRIDHARAN pls interact with the users here , dont look for worked out solution

OpenStudy (kinggeorge):

@waterineyes no.

OpenStudy (anonymous):

See, Let the Dividend be x and the quotient be p in first case when divided by 11 and q in second case when divided by 7.. So, according to Formula I have said above: x = 11p + 9 (11 is Divisor, p is Quotient and 9 is the remainder).. Also, x = 7q + 2 (7 is Divisor, q is quotient and 2 is the remainder)... So, equating the above equations: You get, 11p + 9 = 7q + 2 or, Evaluate q from above: \[q = (11p + 7)/7\] Now all we have to find the least value of p such that the given expression for q becomes a whole number... Suggest any value of p so that (11p + 7) is wholly divisible by 7.. Can you suggest????

terenzreignz (terenzreignz):

Is it 7?

OpenStudy (anonymous):

See, if put p = 0 then it becomes a whole number but it will not constitute to a 3 Digit Number...

terenzreignz (terenzreignz):

So p = 7 that means our x = 11p + 9 is actually x = 86, which indeed, gives a remainder of 2 when divided by 7 and gives a remainder of 9, when divided by 11. But 86 is not a three digit number... ;) Can you do it from here, @PUNYASRIDHARAN ?

OpenStudy (anonymous):

@terenzreignz If we put p = 15 then what is q???

terenzreignz (terenzreignz):

172/7 ?

OpenStudy (anonymous):

No no, If we put p = 14 then what is the value of q?? Can you tell me???

terenzreignz (terenzreignz):

but of course :) it's 163/7

OpenStudy (anonymous):

Check the calculations.. It is 161 not 163 I guess so...

terenzreignz (terenzreignz):

You're right, I'm sorry :) So why did you ask me to calculate?

OpenStudy (anonymous):

I said that the formula is: q = (11p + 7)/7 Use this and get the value of q when p = 14.. Solve and tell me...

terenzreignz (terenzreignz):

You already said it's 161, and I wholly agree (11(14) + 7)/7 (154 + 7)/7 161/7 23

terenzreignz (terenzreignz):

I meant the numerator though.... my bad :)

OpenStudy (anonymous):

So, we are getting q a whole number.. So, p = 14 satisfies the answer... As, x = 11p + 9 from the very first equation of mine: So, put p = 14 here.. x = 11*14 + 9 So, The Number must be 163... So, the three digit number is 163...

terenzreignz (terenzreignz):

Yes, I get that, but that wasn't the question :) The question was how many three digit numbers satisfy the conditions :D

OpenStudy (anonymous):

That is also can be solved very easily: See, here q = (11p + 7)/7 So, if we look at a closer angle to this formula then we arrive on the conclusion that if p is a multiple of 7, then we will get q a whole number.. For example, you first put p = 7 and got 86 but two digit number.. Then I have put put p = 14 (again multiple of 7), we got 163 as number.. Now put 21, 28, 35 etc to find all the three digit numbers... Getting or not??

terenzreignz (terenzreignz):

That, but my idea was get 86 (which is not a thee-digit number) and keep adding 77 which is the least common multiple of 7 and 11 That means adding 77 will not affect the remainder when divided by 7 OR 11. so... 86 (not one of them) 163 (86+77) 240 (163+77) 317 394 471 548 625 702 779 856 933 And that's it, because adding another 77 yields 1010, a four-digit number :) @waterineyes I got you, but I don't want to keep plugging into a linear function :)

OpenStudy (anonymous):

So, the maximum value of p we can plug in to that equation is p = 84 because on putting 91 we are getting 4 Digit Number.. So, we can put 7, 14, 21, 35, 42, 49, 56, 63, 70, 77 and 84.. But at 7, we are getting 86 which is a 2 Digit Number, from above neglect 7.. Now, you will get a 3 Digit Number on 14, 21, 35, 42, 49, 56, 63, 70, 77 and 84.. So, how many numbers will it form: So, there are 10 Three Digit Numbers....

OpenStudy (kinggeorge):

@terenzreignz That's the direct way to do it. @waterineyes there's actually 11 of them

OpenStudy (anonymous):

oh there are 11, then can you please tell me what are the 11 numbers which satisfy the condition given in the question...????

OpenStudy (kinggeorge):

The fastest way to do it, you don't even have to calculate the numbers. Since \[1000=12\cdot77+76\]you know that there are 13 (12+1, since we're counting 0 as well) three digit numbers that satisfy the wanted property less than 1000. Since there are 2 less than 100 (9, 86) there are \[13-2=11\]numbers.

OpenStudy (kinggeorge):

If you want to see all 11 written out, terenz did a good job there 163 240 317 394 471 548 625 702 779 856 933

terenzreignz (terenzreignz):

@KingGeorge Thanks :) I must admit, I've never heard of the Chinese Remainder Theorem until now I'll be sure to look at it and learn it :) Thanks again :D

OpenStudy (kinggeorge):

Just as a warning, it's a little difficult to follow if you look it up on wikipedia.

terenzreignz (terenzreignz):

I know that, but there are good guys out there who actually speak simple English :D I'll seek them out...

OpenStudy (anonymous):

Sorry, I forgot to include 28 in that sequence.. So, total 11 numbers will be there...